Pittsburgh Custom Products (PCP) purchased a new machine for ram-cambering large I beams. PCP expects to bend 85 beams at $2,400 per beam in each of the first 3 years, after which it expects to bend 100 beams per year at $2,900 per beam through year 9. If the company’s minimum attractive rate of return is 15% per year, what is the present worth of the expected revenue?
Answer #1

Revenue in years 1-3 = 85 x 2400 = $ 204000

Revenue in years 4-9 = 100 x 2900 = $ 290000

MARR = 15%

PV = 204000/1.15 + 204000/1.152 + 204000/1.153 + 290000/1.154 + ... + 290000/1.159

PV = $ 1187401.98